The perimeter of a rectangle is 162ft.the ratio of its lenght to width is 5:4.what is the lenght of the rectangle
Answer by Timnewman(323)

let A and B be the lenght and width of the rectangle respectively.
recall that perimeter of a rectangle = 2(A+B)
Now,2(A+B) =162----(1)
also,A/B = 5/4-----(2)
from (1)
2A+2B =162
divide bothside by 2
A+B = 81-----(3)
from (2)
4A = 5B----(4)
from (1)
B = 81-A----(5)
substitute for B in (4)
4A=5(81-A)
Clear bracket,i have
4A=405-5A
add 5A to bothside
4A+5A=405-5A+5A
9A=405
divide bothside by 9
9A/9=405/9
A = 45ft
The lenght of the rectangle is 45ft.